Jill paid $3,000 in interest on her student loans this year. Jill is single, and earned $69,000 from her job as an analyst and $8,500 in interest income. She had no other income this year. What amount of Jill's student loan interest is deductible for AGI this year?

A.$1,500

B. $0

C. $1,250.

D. $2,500
Answer #1

Jill's student loan interest is deductible for AGI this year

D. $2,500

Explanation:-

AGI= 69000+8500-3000= 74,500

MAGI= 74,500+3000= 77,500

Taxpayers who file as single are entitled to a full deduction if their modified adjusted gross income (MAGI) is $65,000 or less, and a partial deduction if their MAGI is over $65,000 but less than $80,000. Anyone with a MAGI over $80,000 can't claim a deduction.

3,000 x (77,500-65,000) / (80,000-65,000)= 2,500